# Right Splitting, Galois Correspondence, Galois Representations and Inverse Galois Problem

### v1

@@ 3.1. Induced Galois representations
From the above theorem, we get coset representatives of $H$, $\{s_{i}\}_{i}$ with $\{\rho(s_{i})\}_{i}$ forming a subgroup with same group structure as $\{s_{i}H\}_{i}=G/H=<sH>$ for some $s\in G$ with $s_{1}\in H$. Hence without loss of generality, let $s_{i}H=s^{i-1}H$ for all $i$. Now let $r_{i}=s_{2}^{i-1}$. Hence $r_{i}H=s_{i}H$. Now since $s_{2}^{n}H=(s_{2}H)^{n}=H$, we have $\tilde{\rho}(s_{2})^{n}=1$. Hence we are done. Other assertions are clear.
@@
Let $H$ be a finite index normal subgroup of $G=Gal(\bar{\mathbb{Q}}/\mathbb{Q})$. Suppose there exists a set of coset representatives of $H$ in $G$ which form a multiplicatively closed subset of $G$. Then if a finite group $M$ is realizable as a Galois group over $\mathbb{Q}$ through a Galois Representation $(\pi,W)$ such that $M\cong\tilde{\pi}(G)=\tilde{\pi}(H)$, then $M\rtimes G/H$ is realisable as a Galois group over $\mathbb{Q}$ (for semidirect product group law as in Thm 3.7).
@@
We can generalize Lemmma  3.4 and Thm 3.7. Consider a closed subgroup $H^{\prime}$ of $G$. We have $H^{\prime}/(H^{\prime}\cap H)\xhookrightarrow{}G/H$.
@@
    $1\rightarrow\rho(H^{\prime}\cap H)\rightarrow\rho(H^{\prime})\rightarrow H^{\prime}/(H^{\prime}\cap H)\rightarrow 1$
@@
    This exact sequence is right split (that is $\rho(H^{\prime})\cong\rho(H^{\prime}\cap H)\rtimes H^{\prime}/(H^{\prime}\cap H)$ for some semidirect product group law) if and only if there exists a set of representatives of $(H^{\prime}\cap H)$-cosets in $H^{\prime}$, $\{s_{1},s_{2},\dots,s_{l}\}$ with $s_{1}\in H^{\prime}\cap H$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(H^{\prime})$.
@@
    Let $H\subset H^{\prime}$. Then the exact sequence is
    $1\rightarrow\rho(H)\rightarrow\rho(H^{\prime})\rightarrow H^{\prime}/H\rightarrow 1$
    If there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ with $s_{1}\in H$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$, then this exact sequence is right split. That is $\rho(H^{\prime})\cong\rho(H)\rtimes H^{\prime}/H$ for some semidirect product group law.
3.2. Galois representations for newforms and Application to Zywina’s case.
Let $p\geq 5$. Fix non-CM newform $f$ of weight $k>1$ on $\Gamma_{1}(N)$. Let $l$ be rational prime as in [8, Zywina]. Let $\pi$ be the Deligne representation associated with $f$ with properties as in 1.1 of [8]. Let $W=E_{\Lambda}^{2}$ and $\pi:G\rightarrow{\rm GL}(W)$ with $\pi(G)\subset{\rm GL}_{2}(\mathcal{O}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow{\rm PGL}_{2}(\mathbb{F}_{\Lambda})$.
Let $p=l$ and assume conditions of [8, 1.2]. Then ${\rm Image}(\tilde{\pi})=\tilde{\pi}(G)={\rm PSL}_{2}(\mathbb{F}_{p})$. The representations $\sigma=\pi|_{H}:H\rightarrow{\rm GL}(W)$ and $\tilde{\sigma}:H\rightarrow{\rm PGL}_{2}(\mathbb{F}_{\Lambda})$ are defined, and ${\rm Image}(\tilde{\sigma})=\tilde{\pi}(H)\subset\tilde{\pi}(G)$.
@@
Let $p\geq 5$ and $H$ be normal in $G$ ($K$ is Galois) and $[G:H]=n<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Then $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
@@
Now, $[{\rm PSL}_{2}(\mathbb{F}_{p}):\tilde{\pi}(H)]=[\tilde{\pi}(G):\tilde{\pi}(H)]\leq[G:H]=n<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Hence $\tilde{\pi}(H)$ is not trivial. Since $H$ is normal in $G$, $\tilde{\pi}(H)$ is normal in $\tilde{\pi}(G)$. Now since ${\rm PSL}_{2}(\mathbb{F}_{p})$ is simple we have that ${\rm PSL}_{2}(\mathbb{F}_{p})=\tilde{\pi}(H)$ otherwise $\tilde{\pi}(H)$ will become non-trivial normal subgroup in ${\rm PSL}_{2}(\mathbb{F}_{p})$.
@@
    For $p\geq 5$, $P{\rm GL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ for some semidirect product group law and the semidirect product is not direct product.
@@
    Let $p\geq 5$. For any semidirect product group law such that $P{\rm GL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$, the semidirect product is not direct product. Furthermore, the automorphism of ${\rm PSL}_{2}(\mathbb{F}_{p})$, given by conjugation by image of generator of $\mathbb{Z}/2\mathbb{Z}$ in $P{\rm GL}_{2}(\mathbb{F}_{q})$, is not inner.
Proof.
(1) From proof of Prop 2.7, we have an exact sequence
$1\rightarrow{\rm PSL}_{2}(\mathbb{F}_{p})\rightarrow{\rm PGL}_{2}(\mathbb{F}_{p})\rightarrow\mathbb{Z}/2\mathbb{Z}\rightarrow 1$
Let $\iota:\mathbb{Z}/2\mathbb{Z}\rightarrow P{\rm GL}_{2}(\mathbb{F}_{p})$ be given by $\iota(\bar{1})=xZ({\rm GL}_{2}(\mathbb{F}_{p}))$ where $x$ is as in Prop 2.7. $\iota$ is a right splitting. Hence above exact sequence is right split. Hence $P{\rm GL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ for some semidirect product group law. Now since, $\begin{bmatrix}1&0\\ 1&1\end{bmatrix}Z({\rm GL}_{2}(\mathbb{F}_{p}))\in PH$ doesn’t commute with $xZ({\rm GL}_{2}(\mathbb{F}_{p}))$, the semidirect product is not direct product.
(2) Let $xZ({\rm GL}_{2}(\mathbb{F}_{p}))$ be the image of generator of $\mathbb{Z}/2\mathbb{Z}$ in $P{\rm GL}_{2}(\mathbb{F}_{q})$ for $x\in{\rm GL}_{2}(\mathbb{F}_{p})$. Then $x\not\in H$, $xZ({\rm GL}_{2}(\mathbb{F}_{p}))\not\in PH$ and $x^{2}\in Z({\rm GL}_{2}(\mathbb{F}_{p}))$. Suppose the semidirect product was direct product. Then for some $\nu_{h}\in Z({\rm GL}_{2}(\mathbb{F}_{p}))$, $xhx^{-1}=h\nu_{h}$ for all $h\in H$. Now, $P{\rm GL}_{2}(\mathbb{F}_{p})=PH\bigsqcup(xZ({\rm GL}_{2}(\mathbb{F}_{p})))PH$. Hence, $xZ({\rm GL}_{2}(\mathbb{F}_{p}))\in Z(P{\rm GL}_{2}(\mathbb{F}_{p}))=\{1\}$. Hence $x\in H$, which gives a contradiction.
Now, suppose automorphism of ${\rm PSL}_{2}(\mathbb{F}_{p})$, given by conjugation by $xZ({\rm GL}_{2}(\mathbb{F}_{p}))$, is inner. Then for some $h^{\prime}\in H$ and $\mu_{h}\in Z({\rm GL}_{2}(\mathbb{F}_{p}))$, $xhx^{-1}=h^{\prime}hh^{\prime-1}\mu_{h}$ for all $h\in H$. Hence $(h^{\prime-1}x)h(h^{\prime-1}x)^{-1}=h\mu_{h}$ for all $h\in H$. Now, $P{\rm GL}_{2}(\mathbb{F}_{p})=PH\bigsqcup((h^{\prime-1}x)Z({\rm GL}_{2}(\mathbb{F}_{p})))PH$. Hence, $(h^{\prime-1}x)Z({\rm GL}_{2}(\mathbb{F}_{p}))\in Z(P{\rm GL}_{2}(\mathbb{F}_{p}))=\{1\}$. Hence $x\in H$, which gives a contradiction.
@@
    For $p\geq 5$, ${\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Thm 3.7).
@@
    The group obtained here is not isomorphic to $P{\rm GL}_{2}(\mathbb{F}_{p})$.
@@
(1) Let $[G:H]=2$. Then $H$ is normal in $G$ ($K$ is Galois). Also $2<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Hence by previous lemma, $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
We could have chosen $K=\mathbb{Q}(i)$ and $s\in G$ as complex conjugation so that $s^{2}=1$ and $\{1,s\}$ become representatives of right cosets of $H$ in $G$. Then by Thm 3.8 we are done.
@@
(4) This follows from (3) and Prop 3.11.
We have results similar to (1), (2) and (3) of Thm 3.12 for certain simple groups of the form ${\rm PSL}_{2}(\mathbb{F}_{q})$.
@@
${\rm PSL}_{2}(\mathbb{F}_{q})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Thm 3.7) for following $q$.
@@
    $q=5^{3},3^{5},3^{4}$ (From 2.2, 2.5, 3 [4] respectively).
Remark 3.13.1.
One has even more general conditions for $q=p^{2}$ (See Thm 3.1 [6]) and $q=p^{4}$ (See 3.3 [6]).
@@ 3.3. Direct sum / tensor product of representations and Galois groups
We have $\pi_{i}:G\rightarrow{\rm GL}(V_{i})$ with $\pi_{i}(G)\subset{\rm GL}_{m_{i}}(\mathcal{O}_{\Lambda})$ and $\tilde{\pi_{i}}:G\rightarrow P{\rm GL}_{m_{i}}(\mathbb{F}_{\Lambda})$ such that $H_{i}\cong{\rm Image}(\tilde{\pi_{i}})$.
Consider Galois representations $(\pi_{i},V_{i}):1\leq i\leq n$ and their direct sum Galois representation $\pi=\bigoplus\limits_{1\leq i\leq n}\pi_{i}:G\rightarrow{\rm GL}\bigoplus\limits_{1\leq i\leq n}V_{i}$.
Now, $\pi(g)={\rm diag}(\pi_{1}(g)\dots\pi_{n}(g))_{m\times m}$ where $m=m_{1}+m_{2}+...+m_{n}$. Hence $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow P{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ with ${\rm Image}(\tilde{\pi})\cong\{(\tilde{\pi}_{1}(g),...,\tilde{\pi}_{n}(g))\ |\ g\in G\}\subset H_{1}\times H_{2}\times\dots\times H_{n}$.
@@
Let $V_{1}$ and $V_{2}$ be finite dimensional vector spaces over a field ${\mathbb{F}}$ with dimensions $m_{1}$ and $m_{2}$ respectively. Given $T_{i}\in{\rm GL}(V_{i})$ for each $i=1,2$, we consider the natural ${\mathbb{F}}$-linear map $(T_{1}\otimes T_{2}):V_{1}\otimes V_{2}\rightarrow V_{1}\otimes V_{2}$ that is also invertible.
Fix bases of $V_{1}$ and $V_{2}$ and the corresponding basis of $V_{1}\otimes V_{2}$. Let $A_{1},A_{2},A$ be the matrices representing $T_{1},T_{2},T_{1}\otimes T_{2}$, respectively with respect to these bases. Then
@@
If the map $T_{1}\otimes T_{2}$ is given by the scalar multiplication by $\lambda\in{\mathbb{F}}^{\times}$, then it follows that $A_{i}=\mu_{i}I$ for some scalars $\mu_{1},\mu_{2}\in{\mathbb{F}}^{\times}$ such that $\lambda=\mu_{1}\mu_{2}$. Thus the map $T_{1}\otimes T_{2}$ descends to a injective group homomorphism
@@
Let ${\mathbb{F}}=E_{\Lambda}$. Then $\Pi(g)=\pi_{1}(g)\otimes\pi_{2}(g)\otimes\dots\otimes\pi_{n}(g)$. Hence $\Pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$ where $m=m_{1}m_{2}\cdots m_{n}$ and $\tilde{\Pi}:G\rightarrow P{\rm GL}_{m}(\mathbb{F}_{\Lambda})$.
Let $F=\mathbb{F}_{\Lambda}$. We observe that $\tilde{\tau}((\tilde{\pi}_{1}(g),\tilde{\pi}_{2}(g)))=\tilde{\pi_{1}(g)\otimes\pi_{2}(g)}=(\tilde{\pi_{1}\otimes\pi_{2}})(g)$. Hence, $\tilde{\tau}(\{(\tilde{\pi}_{1}(g),\tilde{\pi}_{2}(g))\ |\ g\in G\})={\rm Image}(\tilde{\pi_{1}\otimes\pi_{2}})$. Since $\tilde{\tau}$ is injective, we get $\{(\tilde{\pi}_{1}(g),\tilde{\pi}_{2}(g))\ |\ g\in G\}\cong{\rm Image}(\tilde{(\pi_{1}\otimes\pi_{2})})$.
By induction on $n$, we get a well defined injective homomorphism from ${\rm PGL}(V_{1})\times{\rm PGL}(V_{2})\times\dots\times{\rm PGL}(V_{n})$ to ${\rm PGL}(\bigotimes\limits_{1\leq i\leq n}V_{i})$ such that ${\rm Image}(\tilde{\Pi})\cong\{(\tilde{\pi}_{1}(g),...,\tilde{\pi}_{n}(g))\ |\ g\in G\}$.
@@
All the following results for $\rho$ and $\pi$ also hold for $\rho^{\prime}$ and $\pi^{\prime}$ as well as $\tilde{\rho}$ and $\tilde{\pi}$.
@@
    The kernel $ker(\Phi)$ of the projection $\Phi:L\rightarrow\pi(G)$ is isomorphic to $ker(\pi)/(ker(\pi)\cap H)$ and isomorphic to a subgroup of $G/H$.
@@
Let $s_{i}s_{j}H=s_{k}H$ for some $k$, that is $s_{i}s_{j}=hs_{k}$ for some $h\in H$. Now, $\pi(h_{i}s_{i})\pi(h_{j}s_{j})=\pi(h_{i}s_{i}h_{j}s_{j})=\pi(h_{i}h^{\prime}_{j}s_{i}s_{j})$ where $s_{i}h_{j}=h^{\prime}_{j}s_{i}$ since $H$ is normal in $G$. Hence $\pi(h_{i}s_{i})\pi(h_{j}s_{j})=\pi(h_{i}h^{\prime}_{j}hs_{k})$. Let $h_{k}=h_{i}h^{\prime}_{j}h$. Hence $\pi(h_{i}s_{i})\pi(h_{j}s_{j})=\pi(h_{k}s_{k})$. Similarly, $\rho(h_{i}s_{i})\rho(h_{j}s_{j})=\rho(h_{k}s_{k})$.
Hence $\pi(h_{k}s_{k})=I_{m}$ for above $h_{k}$. Let $\rho(h_{k}s_{k})=\theta_{k}$. Then $\theta_{i}\theta_{j}=\theta_{k}$. Hence $\{\theta_{i}\}_{1\leq i\leq n}$ form a group with same group law as $\{(s_{i}H)\}_{1\leq i\leq n}$. Also, for any $i$, $\rho(h_{i}s_{i})=\theta_{i}$ if and only if $\pi(h_{i}s_{i})=I_{m}$. Let $J=\{i\in\{1,2,...,n\}\ |\ ker(\pi)\cap Hs_{i}\neq\emptyset\}$. Hence $ker(\Phi)=\{(I,\theta_{i})\}_{i\in J}$. Thus we have $\Omega:ker(\Phi)\xhookrightarrow{}G/H$ sending $(I,\theta_{i})$ to $s_{i}H$.
Now, since $H\unlhd G$, we have $(ker(\pi)\cap H)\unlhd ker(\pi)$. We also have usual maps $ker(\pi)\xhookrightarrow{}G\rightarrow G/H$. Hence $ker(\pi)/(ker(\pi)\cap H)\xhookrightarrow{}G/H$. Since $G=\bigsqcup_{1\leq i\leq n}Hs_{i}$, we have $ker(\pi)=\bigsqcup_{i\in J}(ker(\pi)\cap Hs_{i})$. Now for $i\in J$ there exists an $h_{i}\in H$ such that $\pi(h_{i}s_{i})=I_{m}$. Hence for $i\in J$, it is easy to see that, $(ker(\pi)\cap Hs_{i})=(ker(\pi)\cap H)(h_{i}s_{i})$. Hence $ker(\pi)=\bigsqcup_{i\in J}(ker(\pi)\cap H)h_{i}s_{i}$. Thus $ker(\pi)/(ker(\pi)\cap H)\cong\{(s_{i}H)\}_{i\in J}\cong ker(\Phi)$.
@@
Now $\tilde{L}/ker(\tilde{\Phi})\cong\tilde{\pi}(G)$ where $\tilde{\Phi}:\tilde{L}\rightarrow\tilde{\pi}(G)$ is the projection map. Since $\tilde{L}$ is finite group, we have $\tilde{L}\cong\tilde{\pi}(G)\iff ker(\tilde{\Phi})$ is trivial $\iff ker(\tilde{\pi})=(ker(\tilde{\pi})\cap H)\iff ker(\tilde{\pi})\subset H$.
We have remarked earlier (3.16) that $ker(\Phi)\xhookrightarrow{}G/H$ via an injective homomorphism say $\Omega$. Let $G^{\prime}$ be the unique subgroup of $G$ such that $H\subset G^{\prime}\subset G$ and $ker(\Phi)=G^{\prime}/H$.
@@
$ker(\Phi)=G^{\prime}/H$ under $\Omega$ $\iff$ $G^{\prime}$ is largest subgroup of $G$ such that $\pi(G^{\prime})=\pi(H)$.

### v2

@@ 3.1. Induced Galois representations and Galois groups
From the above theorem, we get coset representatives of $H$, $\{s_{i}\}_{i}$ with $\{\rho(s_{i})\}_{i}$ forming a subgroup with same group structure as $\{s_{i}H\}_{i}=G/H=<sH>$ for some $s\in G$ with $s_{1}\in H$. Hence without loss of generality, let $s_{i}H=s^{i-1}H$ for all $i$. Now let $r_{i}=s_{2}^{i-1}$. Hence $r_{i}H=s_{i}H$. Now since $s_{2}^{n}H=(s_{2}H)^{n}=H$, we have $\tilde{\rho}(s_{2})^{n}=1$. Hence we are done. Other assertions are clear.∎
@@
Let $H$ be a finite index normal subgroup of $G=Gal(\bar{\mathbb{Q}}/\mathbb{Q})$. Suppose there exists a set of coset representatives of $H$ in $G$ which form a multiplicatively closed subset of $G$. If a finite group $M$ is realizable as a Galois group over $\mathbb{Q}$ through a Galois Representation $(\pi,W)$ such that $M\cong\tilde{\pi}(G)=\tilde{\pi}(H)$, then $M\rtimes G/H$ is realisable as a Galois group over $\mathbb{Q}$ (for semidirect product group law as in Theorem 3.7).
@@
We can generalize Lemma  3.4 and Theorem 3.7. Consider a closed subgroup $H^{\prime}$ of $G$. We have $H^{\prime}/(H^{\prime}\cap H)\xhookrightarrow{}G/H$.
@@
    (3.2) $1\rightarrow\rho(H^{\prime}\cap H)\rightarrow\rho(H^{\prime})\rightarrow H^{\prime}/(H^{\prime}\cap H)\rightarrow 1.$
@@
    The exact sequence 3.2 is right split (that is $\rho(H^{\prime})\cong\rho(H^{\prime}\cap H)\rtimes H^{\prime}/(H^{\prime}\cap H)$ for some semidirect product group law) if and only if there exists a set of representatives of $(H^{\prime}\cap H)$-cosets in $H^{\prime}$, $\{s_{1},s_{2},\dots,s_{l}\}$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(H^{\prime})$.
@@
    Let $H\subset H^{\prime}$. Then the exact sequence 3.2 is 1ρ(H)ρ(H’) H’/H 1. If there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$, then this exact sequence is right split. That is $\rho(H^{\prime})\cong\rho(H)\rtimes H^{\prime}/H$ for some semidirect product group law.
3.2. Galois representations for newforms and application to the cases in work of Zywina
Let $p\geq 5$. Fix non-CM newform $f$ of weight $k>1$ on $\Gamma_{1}(N)$. Let $l$ be rational prime as in [8]. Let $\pi$ be the Deligne representation associated with $f$ with properties as in [8, Section 1.1]. Let $W=E_{\Lambda}^{2}$ and $\pi:G\rightarrow{\rm GL}(W)$ with $\pi(G)\subset{\rm GL}_{2}(\mathcal{O}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow{\rm PGL}_{2}(\mathbb{F}_{\Lambda})$.
Let $p=l$ and assume conditions of [8, Section 1.2]. Then ${\rm Image}(\tilde{\pi})=\tilde{\pi}(G)={\rm PSL}_{2}(\mathbb{F}_{p})$. The representations $\sigma=\pi|_{H}:H\rightarrow{\rm GL}(W)$ and $\tilde{\sigma}:H\rightarrow{\rm PGL}_{2}(\mathbb{F}_{\Lambda})$ are defined, and ${\rm Image}(\tilde{\sigma})=\tilde{\pi}(H)\subset\tilde{\pi}(G)$.
@@ 3.2. Galois representations for newforms and application to the cases in work of Zywina
Let $p\geq 5$ and $H$ be normal in $G$ (i.e. K is Galois over Q) and $G:H]=n<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Then $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
@@
Now, ${\rm PSL}_{2}(\mathbb{F}_{p}):\tilde{\pi}(H)]=[\tilde{\pi}(G):\tilde{\pi}(H)]\leq[G:H]=n<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Hence $\tilde{\pi}(H)$ is not trivial. Since $H$ is normal in $G$, $\tilde{\pi}(H)$ is normal in $\tilde{\pi}(G)$. Now since ${\rm PSL}_{2}(\mathbb{F}_{p})$ is simple we have that ${\rm PSL}_{2}(\mathbb{F}_{p})=\tilde{\pi}(H)$ otherwise $\tilde{\pi}(H)$ will become non-trivial normal subgroup in ${\rm PSL}_{2}(\mathbb{F}_{p})$.∎
@@
    For $p\geq 5$, ${\rm PGL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ for some semidirect product group law and the semidirect product is not direct product.
@@
    Let $p\geq 5$. For any semidirect product group law such that ${\rm PGL}_{2}(\mathbb{F}_{p})\cong{\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$, the semidirect product is not direct product. Furthermore, the automorphism of ${\rm PSL}_{2}(\mathbb{F}_{p})$, given by conjugation by image of generator of $\mathbb{Z}/2\mathbb{Z}$ in ${\rm PGL}_{2}(\mathbb{F}_{p})$, is not inner.
@@
    For $p\geq 5$, ${\rm PSL}_{2}(\mathbb{F}_{p})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Theorem 3.7).
@@
    The group obtained here is not isomorphic to ${\rm PGL}_{2}(\mathbb{F}_{p})$.
@@
(1) Let $[G:H]=2$. Then $H$ is normal in $G$ ($K$ is Galois). Also $2<|{\rm PSL}_{2}(\mathbb{F}_{p})|$. Hence by Lemma 3.10, $\tilde{\pi}(H)={\rm PSL}_{2}(\mathbb{F}_{p})$.
We could have chosen $K=\mathbb{Q}(i)$ and $s\in G$ as complex conjugation so that $s^{2}=1$ and $\{1,s\}$ become representatives of right cosets of $H$ in $G$. Then by Theorem 3.8 we are done.
@@
(4) This follows from (3) and Proposition 3.11.∎
We have results similar to (1), (2) and (3) of Theorem 3.12 for certain simple groups of the form ${\rm PSL}_{2}(\mathbb{F}_{q})$.
@@
The group ${\rm PSL}_{2}(\mathbb{F}_{q})\rtimes\mathbb{Z}/2\mathbb{Z}$ is realisable as Galois Group over $\mathbb{Q}$ (for semidirect product group law in Theorem 3.7) for following $q$.
@@
    $q=5^{3},3^{5},3^{4}$ (From [4, Section 2.2, 2.5, 3]).
Remark $3.1$.
One has even more general conditions for $q=p^{2}$ (See [6, Theorem 3.1]) and $q=p^{4}$ (See [6, 3.3]).
@@ 3.3. Direct sum / tensor product of representations and Galois groups
[…]
@@
Let $V_{1}$ and $V_{2}$ be finite dimensional vector spaces over a general field $\mathbb{F}$} with dimensions $m_{1}$ and $m_{2}$ respectively. Given $T_{i}\in{\rm GL}(V_{i})$ for each $i=1,2$, we consider the natural $\mathbb{F}$}-linear map $(T_{1}\otimes T_{2}):V_{1}\otimes V_{2}\rightarrow V_{1}\otimes V_{2}$ that is also invertible. Fix bases of $V_{1}$ and $V_{2}$ and the corresponding basis of $V_{1}\otimes V_{2}$. Let $A_{1},A_{2},A$ be the matrices representing $T_{1},T_{2},T_{1}\otimes T_{2}$, respectively with respect to these bases. Then
@@
If the map $T_{1}\otimes T_{2}$ is given by the scalar multiplication by $\lambda\in{\mathbb{F}}^{\times}$, then it follows that $A_{i}=\mu_{i}I$ for some scalars $\mu_{1},\mu_{2}\in{\mathbb{F}}^{\times}$ such that $\lambda=\mu_{1}\mu_{2}$. Thus the map $T_{1}\otimes T_{2}$ descends to an injective group homomorphism
@@
We take the special cases ${\mathbb{F}}=E_{\Lambda}$ and then ${\mathbb{F}}=\mathbb{F}_{\Lambda}$ to complete the proof.
Let ${\mathbb{F}}=E_{\Lambda}$. Then $\Pi(g)=\pi_{1}(g)\otimes\pi_{2}(g)\otimes\dots\otimes\pi_{n}(g)$. Hence $\Pi(G)\subset{\rm GL}_{M}(\mathcal{O}_{\Lambda})$ where $M=m_{1}m_{2}\cdots m_{n}$ and $\tilde{\Pi}:G\rightarrow{\rm PGL}_{M}(\mathbb{F}_{\Lambda})$.
[…]
By induction on $n$, we get a well defined injective homomorphism from ${\rm PGL}(V_{1})\times{\rm PGL}(V_{2})\times\dots\times{\rm PGL}(V_{n})$ to ${\rm PGL}(\bigotimes\limits_{1\leq i\leq n}V_{i})$ such that ${\rm Image}(\tilde{\Pi})\cong\{(\tilde{\pi}_{1}(g),...,\tilde{\pi}_{n}(g))\ |\ g\in G\}$.∎
@@
All the following results for the pair ($\rho$, $\pi$) also hold for pairs ($\rho^{\prime}$, $\pi^{\prime}$) as well as ($\tilde{\rho}$, $\tilde{\pi}$).
@@
For $H\unlhd G$ with $[G:H]=n$, we define the projection map $\Psi:L\rightarrow\rho(G)$ by $\Psi((\pi(g),\rho(g)))=\rho(g)$ and the projection map $\Phi:L\rightarrow\pi(G)$ by $\Phi((\pi(g),\rho(g)))=\pi(g)$. Then
@@
    We have $ker(\Phi)\cong ker(\pi)/(ker(\pi)\cap H)$ and it can be seen as a subgroup of $G/H$.
@@
Let $s_{i}s_{j}H=s_{k}H$ for some $k$. Hence for some $h_{k}\in H$, $\pi(h_{i}s_{i})\pi(h_{j}s_{j})=\pi(h_{k}s_{k})$ and $\rho(h_{i}s_{i})\rho(h_{j}s_{j})=\rho(h_{k}s_{k})$. Hence $\pi(h_{k}s_{k})=I_{m}$ for above $h_{k}$. Let $\rho(h_{k}s_{k})=\theta_{k}$. Then $\theta_{i}\theta_{j}=\theta_{k}$. Hence $\{\theta_{i}\}_{1\leq i\leq n}$ form a group with same group law as $\{(s_{i}H)\}_{1\leq i\leq n}$. Also, for any $i$, $\rho(h_{i}s_{i})=\theta_{i}$ if and only if $\pi(h_{i}s_{i})=I_{m}$. Let $J=\{i\in\{1,2,...,n\}\ |\ ker(\pi)\cap Hs_{i}\neq\emptyset\}$. Hence $ker(\Phi)=\{(I,\theta_{i})\}_{i\in J}$. Thus we have $\Omega:ker(\Phi)\xhookrightarrow{}G/H$ sending $(I,\theta_{i})$ to $s_{i}H$.
We have the usual maps $ker(\pi)\xhookrightarrow{}G\rightarrow G/H$. Hence $ker(\pi)/(ker(\pi)\cap H)\xhookrightarrow{}G/H$. Also, $ker(\pi)=\bigsqcup_{i\in J}(ker(\pi)\cap Hs_{i})$. Now for $i\in J$ there exists an $h_{i}\in H$ such that $\pi(h_{i}s_{i})=I_{m}$. Hence for $i\in J$, $(ker(\pi)\cap Hs_{i})=(ker(\pi)\cap H)(h_{i}s_{i})$. Hence $ker(\pi)=\bigsqcup_{i\in J}(ker(\pi)\cap H)h_{i}s_{i}$. Thus
$ker(\pi)/(ker(\pi)\cap H)\cong\{(s_{i}H)\}_{i\in J}\cong ker(\Phi).$
@@
[…]
By Lemma 3.16, $ker(\Phi)$ can be realized as a subgroup of $G/H$ and we denote this embedding by $\Omega:ker(\Phi)\xhookrightarrow{}G/H$. Let $G^{\prime}$ be the unique subgroup of $G$ such that $H\subset G^{\prime}\subset G$ and $ker(\Phi)=G^{\prime}/H$.
@@
$ker(\Phi)=G^{\prime}/H$ under $\Omega$ $\iff$ $G^{\prime}$ is the largest subgroup of $G$ such that $\pi(G^{\prime})=\pi(H)$.